1.     Suppose demand and supply for eggs are given by the following equation:

Qd = 100-20P

Qs = 10 at 40P

            Where Qd = number of eggs purchased yearly

            Where Qs = number of eggs farms would sell yearly

            Where P   = price per dozens of eggs

a)     Fill in the following table

Price Per Dozen

Quantity Demanded (Qd)

Quantity Supplied (Qs)

5.00

6.00

6.50

7.00

7.50

b)     Use the information to find the equilibrium price and equilibrium quantity,

c)     Graph the demand and supply curves and identify the equilibrium price and quantity.


1
Expert's answer
2021-09-09T11:36:01-0400

Given:

The demand function for eggs is: Qd = 100 − 20P

Qd = 100 - 20P

The supply function for eggs is: Qs = 10 + 40P

Qs = 10 + 40P

To Find:

  • The following table with quantity demand and supply information:
  • The equilibrium price and quantity:
  • The graphical representation:

Part-A)

  • The following table with quantity demand and supply information:

The QD will be derived when inserting the price 5:

Qd = 100 − 20P

Qd = 100 − 20 (5)

Qd = 100 − 100

Qd = 0

Qd = 100 - 20PQd = 100 - 20 5Qd = 100 - 100Qd = 0

The QS will be derived when inserting the price 5:

Qs = 10 + 40P

Qs = 10 + 40 (5)

Qs = 10 + 200

Qs = 210

Qs = 10 + 40P

Qs = 10 + 40(5)

Qs = 10 + 200

Qs = 210

Therefore, the table will be filled as shown through above calculations:


Tabular Data:



Thus, in the above-shown table, as demand is less or negative to that of supply, there occurs the situation of EXCESS SUPPLY.


Part-B)


equilibrium is that point where the quantity demand and quantity supply are equal:


Qd=QS

"100-20P=10+40P"

"90=60P"

P"=1.5 (EQUILIBRIUM -PRICE)"

"Q=70 (EQUILIBRIUM -QUANTITY)"


Part-C)